QUESTION 2 Incorrect Mark 0.00 out of 10.00P Flag question Analysis and Interpretation of Profitability Balance sheets and income statements for Target Corporation follow. Income Statement For Fiscal Years Ended ( millions) 2006 2005 2004 Sales Credit card revenues Total revenues Cost of sales Selling, general and administrative expenses Credit card expenses Depreciation and amortization Earnings before interest and income taxes Net interest expense Earnings before income taxes Provisions for income taxes Net earnings 51,271 45,682 40,928 1,349 57 1,097 52,620 46,839 42,025 4,927 31,445 28,389 8,657 722 1,409 1,259 1,098 4,323 3,601 3,159 556 3,860 3,03 2,603 984 $2,408 $1,885 1,619 11,185 9,797 776 737 463 570 1,452 1,146Balance Sheet ($ millions, except footnotes) January 28, 2006 January 29, 2005 Assets Cash and cash equivalents Credit card receivables Inventory Other current assets Total current assets Property and equipment Land Buildings and improvements Fixtures and equipment Computer hardware and software Construction-in-progress Accumulated depreciation Property and equipment, net Other noncurrent assets Total assets Liabilities and shareholders investment Accounts payable Accrued and other current liabilities Current portion of long-term debt and notes payable 1,648 5,666 5,838 1,253 14,405 $ 2,245 5,069 5,384 1,224 13,922 4,449 14,174 3,219 2,214 1,158 (6,176) 19,038 1,552 34,995 3,804 12,518 2,990 1,998 962 (5,412) 16,860 1,511 32,293 6,268 2,567 753 5,779 1,937 5041,552 1,511 Other noncurrent assets Total assets Liabilities and shareholders investment Accounts payable Accrued and other current liabilities Current portion of long-term debt and notes payable Total current liabilities Long-term debt Deferred income taxes Other noncurrent liabilities Shareholders investment Common stock Additional paid-in-capital Retained earnings Accumulated other comprehensive income (loss) Total shareholders investment Total liabilities and shareholders equity $ 34,995 32,293 6,268 2,567 753 9,588 9,119 851 1,232 5,779 1,937 504 8,220 9,034 973 1,037 73 2,121 12,013 74 1,810 11,148 14,205 13,029 $ 34,995 32,293 a. Compute ROE for 2006 Do not round until your final answer. Round answers to two decimal places ROE b. Confirm that ROE equals ROE computed using the component measures for profit margin, asset X%a. Compute ROE for 2006 Do not round until your final answer. Round answers to two decimal places. ROE b. Confirm that ROE equals ROE computed using the component measures for profit margin, asset turnover, and financial leverage using: ROE PM*AT* FL X% Compute the components of ROE Do not round until your final answer. Round answers to two decimal places PM AT- c. Compute adjusted ROA. Assume a tax rate of 38.3% Do not round until your final answer. Round your answer to two decimal places Adjusted ROA X% Check
Answer #1
1)
2006 ROE = Net Margin/ Average Shareholder Equity
Year 2006 2005
Net Margin $       2,408.00
Total shareholders' investment $     14,205.00 $     13,029.00
Average Shareholder Equity = (14205+13029)/2 $     13,617.00
ROE = 2408/13617 17.68%
2)
ROE = profit margin x asset turnover x financial leverage
Profit Margin = Net margin /Total Sales = 2408 /52620 4.58%
Asset Turnover = Total sales/ Average Total assets
Average Total assets = (34995+32293)/2 $     33,644.00
Asset Turnover = 52620/33644 1.56
Financial Leverage = Average Total Assets/ Average Total Shareholder equity
Financial Leverage = 33,644/13617 2.47
ROE = 4.58% x 1.56 x 2.47 17.68%
3)
ROA = Net income + (interest Expense x (1 - tax rate))/ Average Total assets
Adjusted ROA = 2408 + (463 x (1 - 38.3%))/ 33644 8.01%